four times a number increased by 16 is 67. Find the number

Answer by farohw(175)


Let n be a number. Four times a number increased by 16 is 67:

4*n + 16 = 67

solving,

4n + 16 = 67

4n = 67 - 16

4n = 51

4n/4 = 51/4

n = 51/4

Verify:

If n = 51/4 then,

4(51/4) + 16 = 67

51 + 16 = 67

67 = 67